OpenStudy (anonymous):

Find the value of the constant C for which the integral \[\int\limits_{0}^{\infty}(x/x^2+1)−(C/3x+1)dx\] converges. Evaluate the integral for this value of C.

OpenStudy (dumbcow):

ok after integrating \[= (1/2)\ln(x^{2}+1) - (C/3)\ln(3x+1)\] now im having trouble finding the limit as x goes to infinity

OpenStudy (anonymous):

you know to circumvent that indeterminate form, we can combine the fractions

OpenStudy (anonymous):

Like this: ln(x^2 + 1) C ln(3x + 1) ----------- - ------------ = 2 3

OpenStudy (anonymous):

Then : 3 ln(x^2 + 1) - 2C ln(3x + 1) ----------------------------- 6

OpenStudy (anonymous):

which is really: (x^2 + 1)^3 1/6 ln ----------- (3x + 1)^2C

OpenStudy (dumbcow):

ok \[\frac{3\ln(x^{2}+1) - 2C*\ln(3x+1)}{6}\] yes

OpenStudy (anonymous):

right, apply the log properties to what you wrote and you should get what i got: (x^2 + 1)^3 1/6 ln ----------- (3x + 1)^2C

OpenStudy (dumbcow):

yes got that, now find a C that allows this to converge as x->inf

OpenStudy (anonymous):

now for that to converge, we have to have the same powers of x on the top and bottom

OpenStudy (anonymous):

Notice that the top has a polynomial whose degree is 6 and the bottom has a degree of 2c

OpenStudy (dumbcow):

oh right, haha thanks c = 3 converges to ln(1/3)

OpenStudy (anonymous):

to be the same, we simply have to solve this: 2C = 6

OpenStudy (anonymous):

which is c=3

OpenStudy (dumbcow):

well 1/6 * ln(1/3)

OpenStudy (anonymous):

hmm, that was a good problem

OpenStudy (dumbcow):

thanks for the help

OpenStudy (anonymous):

np

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

pretty good for 1:00 in the morning

OpenStudy (dumbcow):

yeah i agree good calc problem...11 pm here so not as impressive

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

were r u at? not North America i gues

OpenStudy (anonymous):

guess*

OpenStudy (dumbcow):

no i am..AZ pacific time

OpenStudy (anonymous):

2 hour time difference approx.